OpenStudy (anonymous):

Determine whether the ratios in each pear are proportioanal. 5/8, 12/20

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ha ha funny

OpenStudy (jamesj):

Two ratios a/b and c/d are proportional if ad = bc. Here you have a/b = 5/8 and c/d = 12/20. i.e., a = 5 b = 8 c = 12 d = 20 That being the case, does ad = bc?

OpenStudy (anonymous):

no

OpenStudy (jamesj):

And btw, this relation ad = bc is true because if \[ \frac{a}{b} = \frac{c}{d} \] then there exists some number, k say, such that \[ \frac{a}{b} = \frac{ka}{kb} \] For example, 1/2 = 2/4 and here k = 2. That being the case c = ka d = kb Thus ad = a(kb) = abk and bc = b(ka) = abk Hence a/b = c/d if and only ad = bc.

OpenStudy (jamesj):

Right, in this case, they are not proportional.

OpenStudy (anonymous):

I like your answers becuase you explain them alot beter than any one else /THANKS!

OpenStudy (jamesj):

thanks

OpenStudy (anonymous):

4/10,2/5 yes or no

OpenStudy (anonymous):

James what grade are you in?

OpenStudy (jamesj):

yes, 4/10 = 2/5. I've finished my studies.

OpenStudy (anonymous):

kewl !

OpenStudy (anonymous):

24/15 ,4/3

OpenStudy (jamesj):

Well, if those two are proportional, you can see that k must be 5 because 15 = 5 x 3 But ...

OpenStudy (anonymous):

what on earth???

OpenStudy (anonymous):

I dont understand

OpenStudy (jamesj):

if 24/15 = 4/3 then there exists some number k such that \[ \frac{24}{15} = \frac{4k}{3k} \] Well, comparing denominators, that means 15 = 3k. Hence k = 5. But if that's right, then 4k = 20. But then the numerators aren't the same. So they are not proportional.

OpenStudy (jamesj):

Another example, is it true that 24/15 = 8/5 ?

OpenStudy (jamesj):

Well, if that is true, there is a number k so that \[ \frac{24}{15} = \frac{8k}{5k} \] that is, 8k = 24 and 5k = 15. Well, k = 3 fits the bill because 8 x 3 = 24 and 5 x 3 = 15. Hence they are proportional, or equal.

OpenStudy (anonymous):

yes